Circular No. 8403

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PKS 0405–385 G. Cimò, S. Ellingsen, S. Carter, and P. McCulloch, University of Tas-

mania; and D. Jauncey and J. Lovell, Australia Telescope National Facility, CSIRO, write: “We have observed a new episode of rapid variability in the flux density of the quasar PKS 0405–385 on hourly time scales. VLBI observations with the Australian Long Baseline Array at 2.2 GHz in Apr. 2004 revealed variations at a level of 10 percent with a period of 2 hr superimposed on a mean flux density of ∼ 1.6 Jy. Observations at 6.7 GHz with the University of Tasmania Ceduna 30-m radio telescope over 10 days in August show that this variability currently continues. Since its discovery in 1996 (IAUC 6418), PKS 0405–385 has shown episodic periods of extremely fast (intra-hour) variability at centimeter radio wavelengths due to interstellar scintillation. The source has alternated between long periods (years) of quiescence and short (a few months) periods of intense activity. The reasons for the sudden change in behavior are not understood, and observations at radio and other wavelengths during the current active phase are encouraged.”

SUPERNOVA 2004ef IN UGC 12158 G. Folatelli, Las Campanas Observatory, together with M. Hamuy and

N. Morrell, on behalf of the Carnegie Supernova Project, report that two optical spectra (range 400–730 nm) of SN 2004ef (cf. IAUC 8399), taken by W. Krzeminski on Sept. 7.16 UT with the Las Campanas 2.5-m du Pont telescope (+ Modular Spectrograph) and by J. Maza and E. Treister on Sept. 8.18 with the 6.5-m Clay telescope (+ Boller & Chivens Spectrograph), reveal that this is a type-Ia supernova, supported by the presence of Si II at 635.5, 413.0, and 597.2 nm and by S II at 545.4 and 564.0 nm. Similar spectral features have been observed in spectra of the normal type-Ia supernovae 1994D and 1999ee at 10 days before maximum, which suggests an early phase for 2004ef. The minimum of the Si II 635.5-nm line yields high expansion velocities of 16200 and 14500 km/s on each night, respectively, assuming the NED host-galaxy recession velocity of 9290 km/s. An upper limit of 0.05 nm is given to the equivalent width of the interstellar Na I D line, indicating a low reddening by dust in the host galaxy.

COMET C/2004 Q1 (TUCKER) J. J. Gonzalez, Leon, Spain, reports visual total mag 12.9 and coma

diameter 1′.5 on Aug. 26.14 UT (0.20-m reflector).

2004 September 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green